OpenStudy (anonymous):

20) What does z equal in the solution of the system of equations below? 6x-7z=13 8y+2z=14 7x+z=6 a)-2 b)2 c)-1 d)1

OpenStudy (amistre64):

x = (6-z)/7 ; Eq3 6x -7z = 13 ; Eq1 6(6-z)/7 -7z = 13 36 -6z -49z = 91 55z = 55 z = 1 ... d

OpenStudy (amistre64):

ack!!..typoed

OpenStudy (amistre64):

-55z = 55 z = -1 ; which is c